. An employee’s weekly income in a certain online job follows a normal distribution with a mean of PHP1,000 and a standard deviation of PHP100. What is the z value for the income of:

1
Expert's answer
2022-06-07T15:31:40-0400

a)


"z=\dfrac{x-\mu}{\sigma}=\dfrac{1100-1000}{100}=1"



b)


"z=\dfrac{x-\mu}{\sigma}=\dfrac{900-1000}{100}=-1"